OpenStudy (anonymous):

Initially, you have lost 20,000 dollars in the stock market and you continue to lose 350 dollars per month. In how many months will it be before your losses total 31600 dollars, thus your balance is - 31600?

OpenStudy (anonymous):

it is not 90

OpenStudy (anonymous):

Let x = number of months Let y = account balance Initially, you lost 20,000, so when x = 0 (in the beginning), y = -20,000. Use y = mx + b Plug in, solve for b. Then use that equation to find what x equals when y = -31600 (x will be the number of months that you are looking for when y, the account balance, equals that).

OpenStudy (anonymous):

okay ill try that

OpenStudy (anonymous):

okay so if x is 0 and im finding y=mx+b, then b would be -20,000 because 350 times 0 would be 0

OpenStudy (anonymous):

Correct. And that makes sense, right? The question said your initial loss was 20,000, so it makes sense that you'd be starting from a deficit of 20,000, yes?

OpenStudy (anonymous):

so would my next equation be set up -316,000=x+(-20,000) since im finding x

OpenStudy (anonymous):

Don't forget your slope. What is m? It's a given, you just have to figure it out from the problem given.

OpenStudy (anonymous):

so my equation would be -316,000=-350x+(-20,000)

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

i get 5920/7

OpenStudy (anonymous):

Watch your y... The problem states that it is -31600, not 316000.

OpenStudy (anonymous):

oh thank you